OpenStudy (nightmarenight):

Solve the equation. 4x2 – 8 = 56

OpenStudy (anonymous):

4x2-8=56+8

OpenStudy (anonymous):

4x=64

OpenStudy (anonymous):

x=16

OpenStudy (doc.brown):

Just to be clear, is the question\[4x^2-8=56\]?

OpenStudy (doc.brown):

If it is\[4x^2-8=56\]Add 8 to both sides\[4x^2-8+8=56+8\]\[4x^2+0=64\]\[4x^2=64\]Divide both sides by 4\[\frac{4x^2}{4}=\frac{64}{4}\]\[\frac{4}{4}x^2=16\]\[1x^2=16\]\[x^2=16\]Find the square root of both sides\[\sqrt{x^2}=\sqrt{16}\]\[x=4\]